OpenStudy (anonymous):

Find the sum of 2 radical 6 and radical 24. A.4 square root 6 B.2 square root 30 C.2 square root 6+4 square root 3 i think its B

OpenStudy (across):

\[2\sqrt{6}+\sqrt{24}\]

OpenStudy (across):

So, why do you think it's B?

OpenStudy (anonymous):

cuz dont you add just the roots

OpenStudy (across):

Do you agree with this:\[\sqrt{4}\sqrt{6}+\sqrt{24}?\]

OpenStudy (across):

What happens when you multiply two radicals?\[\sqrt{4}\sqrt{6}=?\]

OpenStudy (anonymous):

dont they just cancel the root

OpenStudy (across):

\[\sqrt{4}\sqrt{6}=\sqrt{4\cdot6}=\sqrt{24}.\]Do you see that?

OpenStudy (anonymous):

yes i do

OpenStudy (across):

What do we have now?\[\sqrt{24}+\sqrt{24}=2\sqrt{24}.\]Am I correct?

OpenStudy (anonymous):

yes

OpenStudy (across):

What about this?\[2\sqrt{4\cdot6}=2\sqrt{4}\sqrt{6}=2\cdot2\sqrt{6}=4\sqrt{6}.\]Did you follow?

OpenStudy (anonymous):

oh okay i got it know